OpenStudy (anonymous):

Given the equation y − 4 = 3/4 (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

−3/4x + y = 10 3x − 4y = −40 y = 3/4x + 12 y = 3/4x+ 10